# New Permutation Decomposition Techniques For Efficient Homomorphic Permutation

### v5

@@ Abstract.
We start by defining an ideal decomposition form for permutations and propose an algorithm searching depth-1 ideal decompositions. Based on this, we prove the full-depth ideal decomposability of permutations used in specific homomorphic matrix transposition (HMT) and multiplication (HMM) algorithms, allowing them to achieve asymptotic improvement in speed and rotation key reduction. As a demonstration of applicability, substituting the HMM components in the best-known inference framework of encrypted neural networks with our enhanced version shows up to 7.9×7.9$\times$7.9 × reduction in latency.
We further devise a new method for computing arbitrary homomorphic permutations, specifically those with weak structures that cannot be ideally decomposed. We design a network structure that deviates from the conventional scope of decomposition and outperforms the state-of-the-art technique with a speed-up of up to 1.69×1.69$\times$1.69 × under a minimal rotation key requirement.
Homomorphic Encryption, Homomorphic Permutation, Permutation Decomposition
@@ 1. Introduction
Homomorphic permutation often dominates the complexity of homomorphic matrix operations, making its optimization critical for improving the performance of HE-based privacy-preserving applications. Specifically, homomorphic permutation necessitates a substantial number of ciphertext rotations (Halevi and Shoup, 2014, 2018) which are among the most computationally intensive operations in HE, involving multiple number-theoretic transformations over the polynomial ring. Moreover, distinct rotation keys are required to achieve rotations of varying distances, each of which can be significantly larger than the ciphertext itself. The need for numerous rotation keys introduces considerable overhead in both communication and storage (Lee et al., 2023; Chan et al., 2024).
@@
Another compatible optimization is permutation decomposition. It aims to decompose a permutation into multiple simpler ones, each requiring fewer rotations and keys. This concept was first proposed by Gentry et al. (Gentry et al., 2012) where they used the Benes network (Beneš, 1964) for general permutation decomposition. Yet, the application of permutation decomposition in privacy-preserving computations is limited (Meftah et al., 2021, 2022), as existing general methods (Gentry et al., 2012; Halevi and Shoup, 2014) lack adaptability for practical permutations.
@@ 1.1. Technique Overview & Contributions
We begin by analyzing the potential optimal enhancements that decompositions can bring to homomorphic permutations. This leads us to define the ideal decomposition form, which adaptively reduces the complexity of the homomorphic permutation from polynomial to logarithmic, based on the non-zero diagonal distribution of the permutation matrix (Section 2). The rest of the paper revolves around achieving or closely approximating the defined ideal optimization for specific and arbitrary permutations:
@@ 2.1. Notations
We denote scalars and polynomials by lowercase letters, vectors by bold lowercase letters, and matrices by uppercase letters. $A[i,j]$ represents the element in the $i$-th row and $j$-th column of matrix $A$. We also use $A(k,l)_{r}$ to indicate the element in the $k$-th diagonal and $l$-th row of $A$, where $A(k,l)_{r}=A[l,k+i]$. Note that index computation is always implicitly reduced by the matrix’s dimension. Moreover, we denote element-wise product by $\odot$.
@@ 2.2. Batch-encoding Homomorphic Encryption
Typically, for a vector $\mathbf{z}$ to be encrypted, a batch-encoding HE scheme first maps it to some polynomial ring space $R=\mathbb{Z}[X]/(X^{N}+1)$ with $N$ a power of $2$. The mapped plaintext polynomial, denoted $pt(\mathbf{z})$, is then encrypted into a ciphertext $ct(\mathbf{z})$ in $R_{Q_{\ell}}^{2}=(\mathbb{Z}_{Q_{\ell}}[X]/(X^{N}+1))^{2}$ where $Q_{\ell}=\prod_{i=0}^{\ell}q_{i},0\leq\ell<L$ for some pre-determined modulus chain $Q=[q_{0},\dots,q_{L-1}]$. Such ciphertext generally supports the following homomorphic operations (see Appendix A for more details):
@@
    $\texttt{Mult}(ct(\mathbf{z}_{1}),ct(\mathbf{z}_{2}),rlk)$: For the input ciphertexts $ct(\mathbf{z}_{1})\in R_{Q_{\ell}}^{2}$ and $ct(\mathbf{z}_{2})\in R_{Q_{\ell^{\prime}}}^{2}$, compute the tensor product $ct(\mathbf{z}_{1})\otimes ct(\mathbf{z}_{2})\bmod{Q_{\min(\ell,\ell^{\prime})}}$ in $R_{Q_{\min(\ell,\ell^{\prime})}}^{3}$. Then switch it back to $R_{Q_{\min(\ell,\ell^{\prime})}}^{2}$ using a relinearization key $rlk$ and output it as a valid ciphertext of $\mathbf{z}_{1}\odot\mathbf{z}_{2}$.
@@
  1. (1)
    Batch-encoding HE only natively supports homomorphic operations with a limited multiplication depth of $L-1$, as $Q_{L}$ provides at most $L-1$ moduli for Rescale operations.
  2. (2)
[…]
  3. (3)
@@ 2.3. Homomorphic Permutation
Given a ciphertext $ct(\mathbf{x})$ encrypting an $n$-dimensional plaintext vector $\mathbf{x}$ and a length-$n$ permutation $p$ to be applied to $\mathbf{x}$, the homomorphic permutation $ct(p(\mathbf{x}))$ can be represented as a homomorphic linear transformation (HLT) $ct(U\mathbf{x})$, where $U$ is the matrix representation of $p$. $ct(U\mathbf{x})$ can be computed by:
@@ 2.3.1. Permutation Decomposition
Equations 1 and 2 imply that the complexity of homomorphic permutation is closely related to the number of non-zero diagonals in its matrix representation. As mentioned in Section 1, permutation decomposition aims to factorize the permutation into factors possessing sparser non-zero diagonals. This contains two potential benefits: reduced ciphertext rotations and rotation keys.
Gentry et al. (Gentry et al., 2012) utilized the Benes network for permutation decomposition in homomorphic permutation. For any permutation $U$ of length $n$, this approach executes $\log{n}$ rounds of decomposition. In the initial iteration, $U$ is factorized into three components: $U=U^{\sigma}U^{\rho}U^{\tau}$, with $U^{\rho}$ serving as the input of the next round. $U^{\sigma}$ and $U^{\tau}$ contain only non-zero diagonals in {±$\{\pm 2^{\log{n}-i-1},0\}$, where $i$ indicates the iteration round. To our knowledge, the Benes network-based technique (Gentry et al., 2012; Halevi and Shoup, 2014) remains the first and optimal general decomposition method for homomorphic permutation.
Although the Benes network reduces the rotation complexity to roughly $O(4\log{n})$, it fails to adjust complexity based on the distribution of non-zero diagonals, making it counterproductive for highly structured permutations. Another limitation of the Benes network is its significant depth overhead, which can easily exceed the depth supported by homomorphic encryption schemes. While the depth can be reduced by collapsing network levels (Halevi and Shoup, 2014), the complexity of the merged permutations increases as well.
@@
(Ma et al., 2024) Let the $n\times n$ matrix $U$ be the result of multiplying square matrices $U_{L}$ and $U_{R}$ of the same size. Then, $U[i,j]=\sum_{a=0}^{n-1}{{U_{L}[i,a]\cdot U_{R}[a,j]}}$. Denote the diagonal index of $U[i,j]$ as $k=j-i$, the diagonal index of $U_{L}[i,a]$ as $k_{1}=a-i$, and the diagonal index of $U_{R}[a,j]$ as $k_{2}=j-a$. It follows that $k_{1}+k_{2}=k\bmod{n}$.
Subsequent work designing decompositions for specific permutations gives us new insights (Ma et al., 2024). It proposed two-factor decompositions in the form $U=U_{L}U_{R}$, which includes an intriguing pattern (Theorem 1). This implies that two-factor decompositions can be seen as performing subtraction on the diagonal indices. If $U$ has all its non-zero diagonals distributed within the interval $[-r,r]\mod{n}$, and we designate its factor, $U_{R}$, to possess non-zero diagonals exclusively at $\{r_{c},-r_{c},0\}$ ($r_{c}\leq r$), then $U_{L}$ may have its non-zero diagonals confined to the range $[-r^{\prime},r^{\prime}]$ where $r^{\prime}=\max{(r-r_{c},r_{c}/2)}$. Then $U$ can be substituted by $U_{L}$ for further decomposition, as $U_{R}$ maintains a constant sparsity. This idea allows $U$ to be recursively decomposed into a chain:
@@
The non-zero diagonals of $U_{L_{\ell}}$ progressively contract from both sides of the 00-th diagonal, becoming sparser as $\ell$ increases.
In this paper, we specifically define an ideal decomposition form (Definition 2). Its $O(2\log(r/a))$ rotations and $O(\log(r/a))$ depth indicate that $n$ has been decoupled from the complexity, highlighting an enhanced adaptability to highly structured permutations. Such an optimization effect may represent a potential upper bound achievable through decomposition techniques.
@@
Given a square matrix $U$ with non-zero diagonals distributed in interval $[-r,r]$ with a common difference $a\geq 1$, and an algorithm $f$, we say $f$ is an ideal $-\ell$ decomposition ($1\leq\ell<\log{(r/a)}$) of $U$ if $f(U,\ell)=U_{L_{\ell}}U_{R_{\ell}}U_{R_{\ell-1}}\dots U_{R_{1}}$ satisfies:
@@
    The non-zero diagonals of $U_{L_{\ell}}$ are distributed between $[-r^{\prime},r^{\prime}]$, where $r^{\prime}=r-\sum_{i=1}^{\ell}r_{c}^{(i)}$.
If $f$ is an ideal depth-$\log{(r/a)}-1$ decomposition, then we may also call it an ideal full-depth decomposition.
@@ 3. Searching Ideal Depth-1 Decomposition
To find a full-depth ideal (or approximate ideal) decomposition for a given permutation, we can first search for a depth-1 solution and see if some pattern exists for constructing a $-\ell$ solution for $\ell>1$. This section aims to complete the first step, designing an algorithm to search for a depth-$1$ ideal decomposition of any permutation.
@@ 3.1. Basic Decomposition Pattern
To factorize an $n\times n$ permutation matrix $U$ into a product of two factors $U=U_{L}U_{R}$ satisfying the ideal depth-$1$ decomposition of Definition 2, we first observe how non-zero entries on diagonals of $U_{L}$ and $U_{R}$ contribute to those in $U$. For any non-zero entry $U(k,l)_{r}$ in $U$, it is determined by the $k_{1}$-th diagonal in $U_{R}$ and the $k_{2}$-th diagonal in $U_{L}$, where $k_{2}=k-k_{1}\bmod n$ (Theorem 1). This follows that $U_{R}(k_{1},k+l-k_{1})_{r}$ and $U_{L}(k_{2},l)_{r}$ should be set to $1$ to guarantee the correct generation of $U(k,l)_{r}$. We call this decomposing $U(k,l)$ into $k_{1}$-th diagonal and $k_{2}$-th diagonal (see Figure 1). Though entry $U(k,l)$ can be associated with more than two diagonals in general, it will end up with $U_{R}$ and $U_{L}$ not being permutations, which may damage the composability.
In the following context, we use function ConvertR to describe the operation of decomposing any entry $U(k,l)_{r}$ into the $k_{1}$-th diagonal of $U_{R}$:
(4) $\texttt{ConvertR}(k,l,k_{1})=k+l-k_{1}.$
@@
Figure 1. Decompose single entry in $8\times 8$ permutation matrix
@@
[Decomposing single entry in an $8\times 8$ permutation matrix]
During decomposing entries of $U$ into some desired diagonals of $U_{R}$ and $U_{L}$, conflicts (Definition 1) may occur and damage the correctness of factorization. In $U_{R}$, row conflicts may occur, while column conflicts are impossible; conversely, the opposite holds in $U_{L}$. Additionally, since the column index of the entry to be set to $1$ in $U_{L}$ equals $\texttt{ConvertR}(k,l,k_{1})$, a duality exists between the two matrices, indicating that managing either $U_{L}$ or $U_{R}$ is sufficient to make the other matrix undergo the same process.
Definition 0.
We say there is a row conflict (or column conflict) between two non-zero entries in a permutation matrix, if they are in the same row (or column), and there is a conflict between two non-zero diagonals if one or more pairs of conflicting entries exist on these diagonals.
Algorithm 1 Check One Conflicting Row Between the 00-th and ±$\pm r_{c}$-th Diagonals (CheckRow)
@@
4:$r_{c}$: The non-zero diagonal index of $U_{R}$;
5:$a_{0}$: The row with a potential conflict to be resolved;
@@
17:[Try to move $U_{R}(0,a)_{r}$ to the $r_{c}$-th diagonal]
18:$b\leftarrow M_{0}[a],M_{0}[a]\leftarrow\varnothing,\hat{a}\leftarrow(a-r_{c})\bmod{n},solved\leftarrow 0$
19:if $b\geq 0$ and $M_{-r_{c}}[\hat{a}]=\varnothing$ then
@@
21:     $solved\leftarrow\texttt{CheckRow}(M,n,r_{c},\hat{a},Q,i)$triangleright Resolve new conflicts caused by the movement
@@
23:[Otherwise, try to move $U_{R}(0,a)_{r}$ to the $-r_{c}$-th diagonal]
@@
28:         $solved\leftarrow\texttt{CheckRow}(M,n,r_{c},\hat{a},Q,i)$triangleright Resolve new conflicts caused by the movement
@@
34:return $solved$triangleright Return to the previous recursive call

### v7

@@ Abstract.
We start by defining an ideal decomposition form for permutations and propose an algorithm searching for depth-1 ideal decompositions. Based on this, we prove the full-depth ideal decomposability of permutations used in specific homomorphic matrix transposition (HMT) and multiplication (HMM) algorithms, allowing them to achieve asymptotic improvement in speed and rotation key reduction. As a demonstration of applicability, substituting the HMM components in the best-known inference framework of encrypted neural networks with our enhanced version shows up to a $3.9\times$ reduction in latency.
We further devise a new method for computing arbitrary homomorphic permutations, specifically those with weak structures that cannot be ideally decomposed. We design a network structure that deviates from the conventional scope of decomposition and outperforms the state-of-the-art technique with a speed-up of up to $1.69\times$ under a minimal rotation key requirement.
Homomorphic Encryption, Homomorphic Permutation, Permutation Decomposition, Homomorphic Matrix Operation
@@ 1. Introduction
Homomorphic permutation often dominates the complexity of homomorphic matrix operations, making its optimization critical for improving the performance of HE-based privacy-preserving applications. Specifically, homomorphic permutation necessitates a substantial number of ciphertext rotations (Halevi and Shoup, 2014, 2018) which are among the most computationally intensive operations in HE, involving multiple number-theoretic transformations over the polynomial ring. In addition, distinct rotation keys are required to achieve rotations of varying distances, each of which can be significantly larger than the ciphertext itself. The need for numerous rotation keys introduces considerable overhead in both communication and storage (Lee et al., 2023; Chan et al., 2024).
@@
Another compatible optimization is permutation decomposition. It aims to decompose a permutation into multiple simpler ones, each requiring fewer rotations and keys. This notion was first introduced by Gentry et al. (Gentry et al., 2012) where they used the Benes network (Beneš, 1964) for general-purpose permutation decomposition. However, its adoption in privacy-preserving computation remains limited (Meftah et al., 2021), as the complexity bound provided by existing general decomposition techniques (Gentry et al., 2012; Halevi and Shoup, 2014) is only optimized in the permutation length [^2] Throughout this paper, we consistently refer to the length of a permutation as the dimension of the vector upon which the permutation operates.. This bound tends to be loose and insufficiently adaptable to permutations with real-world usages (Jiang et al., 2018; Rizomiliotis and Triakosia, 2022; Meftah et al., 2022; Ma et al., 2024), where the computational complexity is more precisely determined by the number and distribution of diagonals containing non-zero entries in the matrix representation of the permutations (Halevi and Shoup, 2014; Jiang et al., 2018).
@@ 1.1. Technique Overview & Contributions
We begin by analyzing the potential optimal enhancements that decompositions can bring to homomorphic permutations. This motivates us to define an ideal decomposition form, which adaptively reduces the complexity of the homomorphic permutation from polynomial (square root) to logarithmic in the number of non-zero diagonals of the permutation’s matrix representation. The rest of the paper revolves around achieving or closely approximating the defined ideal optimization for specific and arbitrary permutations:
@@ 2.1. Notations
Scalars and polynomials are denoted by lowercase letters (e.g., $a$), vectors by bold lowercase letters (e.g., $\mathbf{a}$), and matrices by uppercase letters (e.g., $A$). $A[i,j]$ represents the element in the $i$-th row and $j$-th column of matrix $A$. We also use $A(k,l)^{*}$ to indicate the element in the $k$-th diagonal and $l$-th row of $A$, where $A(k,l)^{*}=A[l,k+l]$. Index computations are always implicitly reduced by the matrix’s dimension. In addition, we denote element-wise product by $\odot$.
@@ 2.2. Batch-encoding Homomorphic Encryption
Typically, for an $n$-dimensional vector $\mathbf{z}$ to be encrypted, a batch-encoding HE scheme first maps it to some polynomial ring space $R=\mathbb{Z}[X]/(X^{N}+1)$ where $N$ is a power of two and $n\leq N$. The mapped plaintext polynomial, denoted $pt(\mathbf{z})$, is then encrypted into a ciphertext $ct(\mathbf{z})$ in $R_{Q_{\ell}}^{2}=(\mathbb{Z}_{Q_{\ell}}[X]/(X^{N}+1))^{2}$ where $Q_{\ell}=\prod_{i=0}^{\ell}q_{i},0\leq\ell<L$ for some pre-determined modulus chain $Q=[q_{0},\dots,q_{L-1}]$. Such ciphertext generally supports the following homomorphic operations (see Appendix A for more details):
@@
    $\texttt{Mult}(ct(\mathbf{z}_{1}),ct(\mathbf{z}_{2}),rlk)$: For the input ciphertexts $ct(\mathbf{z}_{1})\in R_{Q_{\ell}}^{2}$ and $ct(\mathbf{z}_{2})\in R_{Q_{\ell^{\prime}}}^{2}$, compute the tensor product $ct(\mathbf{z}_{1})\otimes ct(\mathbf{z}_{2})\bmod{Q_{\min(\ell,\ell^{\prime})}}$ in $R_{Q_{\min(\ell,\ell^{\prime})}}^{3}$. Then switch it back to $R_{Q_{\min(\ell,\ell^{\prime})}}^{2}$ using a relinearization key rlkrlk and output it as a valid ciphertext of $\mathbf{z}_{1}\odot\mathbf{z}_{2}$.
@@
  • $\bullet$
    Batch-encoding HE only natively supports homomorphic operations with a limited multiplication depth of L1L-1, as $Q_{L}$ provides at most L1L-1 moduli for Rescale operations.
  • $\bullet$
    The complexity order of the routines is: $\texttt{Mult}\approx\texttt{Rot}>\texttt{Rescale}>\texttt{CMult}>\texttt{Add}$. Here, Add and CMult are typically performed in the NTT domain, with an asymptotic complexity of $O(N)$. While others require NTT-domain switches with an $O(N\log{N})$ complexity.
  • $\bullet$
@@ 2.3. Homomorphic Permutation
Given a ciphertext $ct(\mathbf{x})$ encrypting an $n$-dimensional plaintext vector $\mathbf{x}$ and a length-$n$ permutation $p$ to be applied to $\mathbf{x}$, the homomorphic permutation $ct(p(\mathbf{x}))$ can be represented as a homomorphic linear transformation (HLT) $ct(U\mathbf{x})$, where $U$ is the matrix representation of $p$. Specifically, $U[i,j]=1$ if $p$ maps the $i$-th entry of $\mathbf{x}$ to position $j$, otherwise $U[i,j]=0$. $ct(U\mathbf{x})$ is computed by:
@@
Figure 1. Depth-2 ideal decomposition on a $32\times 32$ permutation with non-zero diagonal indices in $\{2\cdot i\mid-6\leq i\leq 6\}$
@@ 2.3.1. Permutation Decomposition
Equations 1 and 2 imply that the complexity of homomorphic permutation is closely related to the number of non-zero diagonals in its matrix representation. As mentioned in Section 1, permutation decomposition aims to factorize the permutation into factors possessing fewer non-zero diagonals. This contains two potential benefits: reduced ciphertext rotations and rotation keys.
Gentry et al. (Gentry et al., 2012) utilized the Benes network for permutation decomposition in homomorphic permutation. For any permutation $U$ of length $n$, this approach executes $\log{n}$ rounds of decomposition. In the initial iteration, $U$ is factorized into three components: $U=U^{\sigma}U^{\rho}U^{\tau}$, with $U^{\rho}$ serving as the input of the next round. $U^{\sigma}$ and $U^{\tau}$ contain only non-zero diagonals in $\{\pm 2^{\log{n}-i-1},0\}$, where $i$ indicates the iteration round. To our knowledge, the Benes network-based technique (Gentry et al., 2012; Halevi and Shoup, 2014) remains the first and optimal general decomposition method for homomorphic permutation.
Although the Benes network lowers the rotation complexity to roughly $4\log n$, it fails to tailor this cost to the distribution of non‑zero diagonals, which makes it counter‑productive for highly structured permutations. Specifically, it does not consider the arithmetic sequence among the non-zero diagonal indices and may disrupt this regularity during the decomposition rounds. As a result, the expense of evaluating the decomposed factors with HLT is no longer reduced by the length of the arithmetic sequence and can even surpass the cost of executing the original permutation. Another limitation of the Benes network lies in its significant depth overhead ($2$ additional levels per round), which can easily exceed the depth supported by homomorphic encryption schemes. While the depth can be reduced by collapsing network levels (Halevi and Shoup, 2014), the complexity of the merged permutations may increase as well.
@@
(Ma et al., 2024) Let $U$ be an $n\times n$ matrix formed by the product of two square matrices $U_{L}$ and $U_{R}$. Then, $U[i,j]=\sum_{a=0}^{n-1}U_{L}[i,a]\cdot U_{R}[a,j]$, and the diagonal index k=jik=j-i of $U[i,j]$ can be decomposed as the sum of the diagonal index $k_{L}=a-i$ of $U_{L}[i,a]$ and the diagonal index $k_{R}=j-a$ of $U_{R}[a,j]$, that is, $k=k_{L}+k_{R}\mod n$.
Subsequent work designing decompositions for specific permutations gives us new insights (Ma et al., 2024). It proposed two-factor decompositions in the form $U=U_{L}U_{R}$, which includes an intriguing pattern (Theorem 1). This implies that two-factor decompositions can be seen as performing subtraction on the diagonal indices. If $U$ has all its non-zero diagonals distributed within the interval $r,r]\mod{n}$, and we designate its factor, $U_{R}$, to possess non-zero diagonals exclusively at position $\{r_{c},-r_{c},0\}$ ($r_{c}\leq r$), then $U_{L}$ may have its non-zero diagonals’ indices confined to the range $r^{\prime},r^{\prime}]$ where $r^{\prime}=\max{(r-r_{c},r_{c}/2)}$. Then $U$ can be substituted by $U_{L}$ for further decomposition, as $U_{R}$ maintains a small constant number of non-zero diagonals. This idea allows $U$ to be recursively decomposed into a chain:
@@
The non-zero diagonals of $U_{L_{\ell}}$ progressively contract from both sides of the 0-th diagonal, with their number decreasing as $\ell$ increases.
In this paper, we specifically define an ideal decomposition form (Definition 2). “Ideal” denotes that each round of decomposition halves the number of non‑zero diagonals in the factor decomposed, up to an additive constant of $2$ (excluding the 0-th diagonal). Figure 1 illustrates a depth-$2$ ideal decomposition on a $32\times 32$ permutation matrix whose non-zero diagonal indices follow the arithmetic sequence $s=\{a\cdot i\mid-r/a\leq i\leq r/a\}$ with common difference $a=2$ and distribution range $r=12$. One can observe that the number of non-zero diagonals in $U$ and $U_{L_{1}}$ are halved in their left factors, while the right factors preserve a constant number of diagonals.
A full-depth ideal decomposition requires $O(\log(r/a))$ rotations and a multiplication depth of $\lfloor\log(r/a)\rfloor+1$. Figure 1 shows exactly a full-depth decomposition, where the sub-permutations $U_{R_{1}},U_{R_{2}},U_{L_{2}}$ can be sequentially applied to any input ciphertext $ct(\mathbf{x})$ to compute $ct(U\mathbf{x})$, each requiring a single HLT. This results in a total of $2(\lfloor\log{(r/a)}\rfloor+1)=6$ rotations. Furthermore, if the decomposition is partial with a depth of $\ell$, the $U_{L_{\ell}}$ in the decomposition chain will benefit from a BSGS-based HLT, yielding an overall cost of $O\left(\sqrt{(r/a)/2^{\ell}}+\ell\right)$, due to the preserved arithmetic sequence in $U_{L_{\ell}}$. In contrast to the Benes network, this decomposition scheme decouples the complexity from $n$, implying an enhanced adaptability to structured permutations in which the non-zero diagonals are regularly and sparsely distributed across the permutation matrix. Such an optimization effect may represent a potential upper bound in speed-up achievable through decomposition techniques.
@@
Given a square matrix $U$ with its non-zero diagonals distributed in interval [r,r][-r,r] as an arithmetic sequence $s=\{a\cdot i\mid-r/a\leq i\leq r/a\}$ with the common difference $a\geq 1$, and a decomposition $f$, we say $f$ is an ideal depth-$\ell$ decomposition ($1\leq\ell\leq\lfloor\log{(r/a)}\rfloor$) of $U$ if $f(U,\ell)=\{U_{L_{\ell}},U_{R_{\ell}},U_{R_{\ell-1}},\dots,U_{R_{1}}\}$ satisfies:
  1. (1)
    $U=U_{L_{\ell}}U_{R_{\ell}}U_{R_{\ell-1}}\dots U_{R_{1}}$.
@@
    The non-zero diagonals of $U_{L_{\ell}}$ are distributed between $r^{\prime},r^{\prime}]$, where $r^{\prime}=r-\sum_{i=1}^{\ell}r_{c}^{(i)}$.
If $f$ is an ideal depth-$\lfloor\log{(r/a)}\rfloor$ decomposition, then we may also call it an ideal full-depth decomposition.
@@ 3. Searching Ideal Depth-1 Decomposition
To find a full-depth ideal decomposition for a given permutation, we can first construct a method searching for a depth-$1$ solution. This initial search can then be recursively applied to the resulting factors to achieve further decomposition. Moreover, the depth‑$1$ decomposition factors themselves may reveal interpretable structural patterns that serve as evidence of full-depth decomposability. This section aims to complete the first step, designing a depth-$1$ ideal decomposition search for any permutation.
@@ 3.1. Basic Decomposition Pattern
To factorize an $n\times n$ permutation matrix $U$ into a product of two factors ($U=U_{L}U_{R}$) satisfying the ideal depth-$1$ decomposition of Definition 2, we first observe how non-zero entries on diagonals of $U_{L}$ and $U_{R}$ contribute to those in $U$. For any non-zero entry $U(k,l)^{*}$ in $U$, it is determined by the $k_{R}$-th diagonal in $U_{R}$ and the $k_{L}$-th diagonal in $U_{L}$, where $k_{L}=k-k_{R}\bmod n$ (Theorem 1). This follows that $U_{R}(k_{R},k+l-k_{R})^{*}$ and $U_{L}(k_{L},l)^{*}$ should be set to $1$ to guarantee the correct generation of $U(k,l)^{*}$. We call this decomposing $U(k,l)$ into $k_{R}$-th diagonal and $k_{L}$-th diagonal (see Figure 2). Though entry $U(k,l)$ can be associated with more than two diagonals in general, it will end up with $U_{R}$ and $U_{L}$ not being permutations, which may damage the composability.
In the following context, we use function ConvertR to describe the operation of decomposing any entry $U(k,l)^{*}$ into the $k_{R}$-th diagonal of $U_{R}$:
(4) $\texttt{ConvertR}(k,l,k_{R})=k+l-k_{R}.$
@@
During decomposing entries of $U$ into some desired diagonals of $U_{R}$ and $U_{L}$, conflicts may occur and damage the correctness of factorization. Specifically, we say there is a row (or column) conflict in the matrix if there is more than one non-zero entry in a row (or column). In $U_{R}$, row conflicts may occur, while column conflicts are impossible; conversely, the opposite holds in $U_{L}$. Additionally, since the column index of the entry to be set to $1$ in $U_{L}$ equals $\texttt{ConvertR}(k,l,k_{R})$, a duality exists between the two matrices, indicating that managing either $U_{L}$ or $U_{R}$ is sufficient to make the other matrix undergo a mirroring process.
Figure 2. Decompose single entry in $8\times 8$ permutation matrix: Decomposing single entry in an $8\times 8$ permutation matrix
Algorithm 1 Depth-11 Ideal Decomposition Search
1:
2:$U$: Permutation Matrix of size $n\times n$;
3:$a$: Common difference of the non-zero diagonal indices in $U$;
4:$r$: Distribution bound of the non-zero diagonal indices in $U$.
5:
@@
7:[Step 1: Decompose entries]
[…]
11:for $k$ traversing indices of non-zero diagonals in $U$ do
12:  for $l$ traversing row indices of non-zero entries in the $k$-th diagonal do
13:   if $M_{-r_{c}}[\texttt{ConvertR}(k,l,r_{c})]=\varnothing$ and $k>=r_{c}$ then
14:     $M_{r_{c}}[\texttt{ConvertR}(k,l,r_{c})]\leftarrow k$
15:   else if $M_{r_{c}}[\texttt{ConvertR}(k,l,-r_{c})]=\varnothing$ and $k<=r_{c}$ then
16:     $M_{-r_{c}}[\texttt{ConvertR}(k,l,-r_{c})]\leftarrow k$
17:   else if $r_{c}<k<r_{c}$ then
18:     $M_{0}[\texttt{ConvertR}(k,l,0)\leftarrow k$
19:   else
20:     return $\varnothing,\varnothing$ $\triangleright$ Decomposition not exists
21:   end if
22:  end for
23:end for
24:[Step 2: Resolve conflicts]
[…]
26:for $a$ traversing the keys in $M_{0}$ do
27:  if $M_{-r_{c}}[a]\not$$\varnothing$ or $M_{r_{c}}[a]\not$$\varnothing$ then
28:   Append $a$ to $Q$ $\triangleright$ Add one conflict to list
29:  end if
30:end for
[…]
32:if $|Q|\not=0$ then
33:  $solved\leftarrow\texttt{CheckRow}(M,n,r_{c},Q[0],Q,0)$
34:end if
35:[Step 3: Reconstruct $U_{L}$]
36:if $solved=0$ then
37:  return $\varnothing,\varnothing$ $\triangleright$ Decomposition not exists
38:end if
39:Construct $U_{R}$ from $M$, Compute $U_{L}$ such that $U=U_{L}U_{R}$
40:return $U_{L},U_{R}$ $\triangleright$ Decomposition exists
Algorithm 2 Check One Conflicting Row Between the 0-th and $\pm r_{c}$-th Diagonals (CheckRow)
@@
[…]
@@
17:[Try to move $U_{R}(0,a)^{*}$ to the $r_{c}$-th diagonal]
[…]
19:if $b\geq 0$ and $M_{-r_{c}}[\hat{a}$$\varnothing$ then
@@
21:  $solved\leftarrow\texttt{CheckRow}(M,n,r_{c},\hat{a},Q,i)$ $\triangleright$ Resolve new conflicts caused by the movement
@@
23:[Otherwise, try to move $U_{R}(0,a)^{*}$ to the $r_{c}-$th diagonal]
@@
28:   $solved\leftarrow\texttt{CheckRow}(M,n,r_{c},\hat{a},Q,i)$ $\triangleright$ Resolve new conflicts caused by the movement
@@
34:return solvedsolved $\triangleright$ Return to the previous recursive call